why aren't the magnification of both ocular lenses of a binocular microscope used to calculate total magnification?
Answer #1

1. Why aren't the magnifications of both ocular lenses of binocular microscope used to calculate total magnification?

Because the image only goes through one ocular to reach both eyes. And the lenses are side by side and not in a line with each other.

2.

MT = 15 * 4 = 60x
MT = 15 * 10 = 150x
MT = 15 * 45 = 675x
MT = 15 * 97 = 1455x

3. Limit if resolution (D) = wavelength / (NA condenser + NA objective)

A small wavelength light will allow for a smaller limit of resolution value.

limit the resolution refers to the distance apart two objects must be for the mucroscope to distinguish them as separate objects. smaller D value means better resolution, thus better clarity.

4. Because the shorter the wavelength is the more detail it will show in ocular lens.
Answer #2
Since each lens is capable of magnifying the vision of one eye, thus the magnification is calculated using one lens only
Answer #3
These lenses are parallel, not in series. The light that passes through the sample is magnified by the objective and then the ocular, not objective and then through two oculars. If they were stacked on top of one another, then yes, you would have to take that into account. Close one eye and observe the apparent size of the image...then do it with the other eye. The image stays the same, right! So only the path of light is important. Simply multiply the ocular (10x) by the objective to get the total magnification.
please rate this
Answer #4
These lenses are parallel, not in series. The light that passes through the sample is magnified by the objective and then the ocular, not objective and then through two oculars. If they were stacked on top of one another, then yes, you would have to take that into account. Close one eye and observe the apparent size of the image...then do it with the other eye. The image stays the same, right! So only the path of light is important. Simply multiply the ocular (10x) by the objective to get the total magnification.